OpenStudy (ragingsquirrel):

A rectangular prism with a height of 6 m, a length of 5 m, and a width of 4 m will have a different surface area than a rectangular prism with a height of 5 m, a width of 6 m, and a length of 4 m. True False

OpenStudy (ragingsquirrel):

True, right?

hero (hero):

\[A_{rectangular prism} = 2(ab + bc + ac)\]

OpenStudy (hoblos):

same dimensions ... then it has same area false

hero (hero):

Hoblos, do you want me to do the math for you?

OpenStudy (hoblos):

yes please...

OpenStudy (ragingsquirrel):

...so its true?

OpenStudy (hoblos):

i still go with FALSE

hero (hero):

okay, for the first prism, let a = 4, b = 5, c = 6, then: A = 2(ab + bc + ac) = 2(4*5 + 5*6 + 4*6) = 2(20 + 30 + 24) = 2(50 + 24) = 2(74) = 148 For the second prism, let a = 6, b = 4, and c = 5, then: A = 2(ab + bc + ac) = 2(6*4 + 4*5 + 6*5) = 2(24 + 20 + 30) = 2(24 + 50) = 2(74) = 148 Any questions?

hero (hero):

@hoblos please do not disagree with anything I say without proof.

OpenStudy (hoblos):

so they are not DIFFERENT

hero (hero):

I mis-read the question, my answer was that they are the same and not different. I apologize.

hero (hero):

Anyways, slight oversight by me. I was convinced that the question was asking if they had the same area. I didn't even see "different"

hero (hero):

@RagingSquirrel

OpenStudy (hoblos):

dont worry Hero it was just a little mistake you explained your point well though XD

hero (hero):

I doubt if Raging Squirrel misread the question like I did. He probably honestly believed that they should be different, but I know better.

OpenStudy (ragingsquirrel):

I did actually misread that... thanks for helping guys.

hero (hero):

Next time, use the formula and prove it to yourself, that way I don't have to mis-read it with you.

OpenStudy (ragingsquirrel):

haha, will do.